OpenStudy (anonymous):

Write an equation in standard form of the parabola that has the same shape as the graph of f(x)=-5x^2 , but with vertex at (-16,-11). the equation is y=?

OpenStudy (zale101):

use y = a(x-h)^2 + k to set plug in the numbers of the given vertex, h=x and k=y

OpenStudy (zale101):

so, let(x) be y y = -5x^2 turn it into this form y = a(x-h)^2 + k y = -5(x - (-16))^2 + (-11) y = -5(x +16)^2 -11